Question #4: IS-LM Model: Change in Fiscal Policy (a) Suppose Congress had announced that they were going to increase governmQuestion #4: IS-LM Model: Change in Fiscal Policy (a) Suppose Congress had announced that they were going to increase government spending to G = 400. Assume that (M/P)Sreturns to 1600. Now the set of equations are the following: C = 200 + 0.25YDI = 150 + 0.25Y –1000i T = 200 G = 400(M/P)S= 1600(M/P)d= 2Y –8000i Calculate the new level of equilibrium interest rate (i) and equilibrium output (Y).(b) Calculate the new levels of consumption (C) and investment (I) when government spending increases.(c) Graphically illustratethe change in the IS-LM Model as result of the change in government spending.In your graph draw the initial equilibrium (similar to what you did for Question 3(a)). Clearly indicate the initial equilibrium interest rate and output. Label this point as Point A. Illustrate thenew equilibrium. Be sure to clearly indicate the new levels of equilibrium interest rate and output. Label this point as Point B.
Answer #1

wton 4 Question 4 suppose a suppose Congress had announced that they wwe going to increase government spending to a=400. andNow Equation of LM (AUL! - ri d = ( 4 ) 2) 27-8000i = 1600 2 мою, Рум 24 би di) wwwrkt- 2y = 1600 + 8020 i 21 y (1600 + 80001New level of Consumption 2 + 0,8 ( y -T) 2) C = 200 +0.25 ( 1200-200) les 200 + (0.25 X lovo) = 450 c) es 200 td and Investme